OpenStudy (anonymous):

60% of a number is 90. What is 50% of that same number? A. 30 B. 45 C. 75 D. 150

OpenStudy (anonymous):

its 75

OpenStudy (lucaz):

60%=60/100=0.6 0.6*x=90 x=90/0.6 x=150 -> 60% of 150 = 0.6*150 = 90

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

*poop*

OpenStudy (anonymous):

lolz @planslam